# FreshStack: Building Realistic Benchmarks for Evaluating Retrieval on Technical Documents

### v1

@@ 3.4. Retrieval: Oracle & Inference Setting
A RAG evaluation dataset requires questions, answers, and a corpus with documents, which helps support facts in the answer. In this stage, we retrieve a list of highly relevant unjudged documents from the corpus and construct judgment pools. Since, we are constructing an evaluation dataset and we have curated answers for questions, we retrieve a list of top-k documents using two methods: (1) Inference, relying only on the question and automatic approaches, and (2) Oracle, relying on the gold answer or list of nuggets, to pool diverse documents for relevance judgment in the next stage.
Retrieval Settings. We experiment with multiple settings to increase diversity in our judgment pools. First, we experiment with two techniques in the inference setting:
  • GPT-4o Sub-Questions. Decomposing the original question, we generate a few synthetic sub-questions using GPT-4o, similar to Rosset et al. (2024), concatenated together to retrieve documents.
  • GPT-4o Closed Book Answer. We generate a closed-book answer for the original question with GPT-4o, similar to HyDE (Gao et al., 2023a), and use the closed-book answer to retrieve documents.
Next, in the Oracle setting, we also experiment with two techniques:
  • Stack Overflow Answer: We use the curated Stack Overflow answer as the question to retrieve documents.
  • Stack Overflow Nuggets: We use the list of GPT-4o generated nuggets (Section 3.3), concatenated as the question to retrieve documents.
Retrieval Models. We experiment with five different code and text-aware retrieval models:
  1. (1)
    BM25, a strong lexical baseline in BEIR (Thakur et al., 2021). We utilized the default BM25 implementation available in Pyserini (Lin et al., 2021).
  2. (2)
    BGE (Gemma-2) (Chen et al., 2024) a dense retriever model [^4] BGE Gemma-2: https://huggingface.co/BAAI/bge-multilingual-gemma2 fine-tuned on the backbone architecture of Gemma-2 (9B) (Rivière et al., 2024) with an embedding size of 3584 and 8K context length.
  3. (3)
    E5 Mistral (7B) (Wang et al., 2024b) is a dense retriever model [^5] E5 Mistral 7B: https://huggingface.co/intfloat/e5-mistral-7b-instruct based fine-tuned on the backbone of Mistral 7B (Jiang et al., 2023) with 32 layers and embedding size of 4096.
  4. (4)
    Voyage-large-2666Voyage-large-2: https://docs.voyageai.com/docs/embeddings is a proprietary and general-purpose embedding model optimized for retrieval quality, with a context length of 16K tokens and embedding size of 1536.
  5. (5)
    Fusion, a hybrid retrieval strategy with the four individual models, by normalizing and summing up the top 100 documents and their scores from each model.
3.5. Nugget-Level Support Assessment
@@
This is the final stage in the framework. Traditionally, relevance judgments are conducted on selected pools of retrieved documents, i.e., a human assessor judges the relevance of the question and each provided document pair. Due to computational costs, recent studies experiment with an LLM-as-a-judge (instead of a human assessor) for conducting relevance judgments in information retrieval (Faggioli et al., 2023; Thomas et al., 2024; Upadhyay et al., 2024b, a; Rahmani et al., 2024). Questions in existing IR datasets are traditionally short, making it easier to assess question-document relevance. In contrast, questions in the FreshStack dataset are long and elaborate (between 350–500 tokens in length), containing a mixture of text, code snippets, or outputs, making it challenging to judge question-document relevance directly (Damessie et al., 2016). For instance, retrieved documents may answer a major problem presented in the question, address only part of the question, or contain relevant references and examples, and we need to translate this into a relevance score.
@@
#Queries #Docs #GitHub Avg. N/Q Query Answer Query Answer Rel. Docs/N Rel. Docs/Q
@@
Table 3. FreshStack dataset statistics; Dataset count measures the number of queries, documents, GitHub repositories, and average nuggets per query; Avg. length measures the average text lengths (in tokens); % containing code measures the fraction of queries and answers with code snippets; Relevance judgments measure relevant documents per nugget and query.
@@
Nugget-level Support. Instead of relying on traditional relevance assessments, we simplify the judgment procedure for the LLM and evaluate whether a document supports information (or contains) provided by a nugget. A reminder that a nugget highlights an essential fact of the Stack Overflow question or answer. Judging document relevance at a nugget level is effective as nuggets are factual and short information snippets, reducing the ambiguity often seen during traditional relevance judgments. As we have $n$ nuggets per question, repeating the assessment of $k$ documents for every nugget is prohibitive ($n\times k$). To reduce computational costs, we evaluate top-k documents (a maximum $k=20$) together with the list of all nuggets for a question in a single inference call ($n+k$). We evaluate support judgment with GPT-4o using chain-of-thought prompting (Wei et al., 2022).
4. Dataset Statistics & Evaluation
@@
Upon completion of previous stages, we employ two post-processing steps to ensure high-quality question and answer pairs remain in the dataset, reducing the overall dataset size. In the first post-processing stage, we remove unsupported questions, i.e., questions that do not contain even a single relevant document; this removes, on average, 11.8% of the total questions. [^7] Future work may include these questions as they are potentially valuable to answer, and better retrieval systems may be able to find relevant documents. In the next stage, we aggressively filter by removing questions containing at least one unsupported nugget, i.e., a nugget that is not supported by any documents, reducing on average 34.2% of the total questions.
4.1. Dataset Statistics
@@
FreshStack datasets covers five commonly used topics for programmers: machine learning, computer vision, backend, front-end, and game development, all listed in Table 3. Stack Overflow topics such as LangChain were introduced in 2023, whereas others, like Laravel or Angular, have questions about the latest versions (e.g., Laravel 10, Laravel 11). Each topic has at least 50 questions, and the corpus has at least 25K documents sourced from 4–10 GitHub repositories. The questions are long, containing 350–500 tokens (computed using GPT-4o tokenizer), and at least 50% of the questions and answers contain code snippets. GPT-4o generates around 3–4 nuggets for each question. Each nugget is matched to at least 3 relevant documents with support judgment from GPT-4o, resulting in 5–6 relevant documents per question, across all topics.
4.2. Instance Description
Each FreshStack dataset instance contains the following four components, as shown in Figure 1:
  • Relevance Judgments: Unlike traditional IR benchmarks, such as BEIR (Thakur et al., 2021), which contain question and document-level relevance judgments, FreshStack datasets contain nugget-level relevance judgments for document chunks.
Method Model LangChain (ML) Yolo v7 & v8 (Vision) Laravel 10 & 11 (Backend) Angular 16, 17 & 18 (Front-end) Godot4 (Game Dev)
$\alpha$-N@10 Cov@20 Rec@50 $\alpha$-N@10 Cov@20 Rec@50 $\alpha$-N@10 Cov@20 Rec@50 $\alpha$-N@10 Cov@20 Rec@50 $\alpha$-N@10 Cov@20 Rec@50
@@ 4.2. Instance Description
Table 4. Pooling results achieved by retrieval baselines (including fusion) in inference or oracle settings for constructing the FreshStack dataset. $\alpha$-N@10 denotes $\alpha$-nDCG@10, Cov@20 denotes Coverage@20 and Rec@50 denotes Recall@50. Stack Overflow Answer & Answer Nuggets methods rely on the gold answer for retrieval (oracle setting), whereas, other methods do not rely on the gold answer for retrieval. Overall, we highlight the best result in bold for each setting.
4.3. Retrieval Evaluation Metrics
@@
Information retrieval (IR) evaluation follows the Cranfield paradigm (Voorhees, 2009), focusing on individual document relevance, independent of other documents. This is used to construct standard test collections, such as BEIR (Thakur et al., 2021) and TREC datasets such as the Deep Learning (DL) track (Craswell et al., 2022, 2023, 2024). However, diversity in search (Carbonell and Goldstein, 1998; Santos et al., 2015; Wu et al., 2024) penalizes information redundancy within retrieved documents to enrich information content and improve efficiency. Therefore, we evaluate retrieval systems using three diversity-focused metrics: $\alpha$-nDCG@10 for diversity and relevance, Coverage@20 for nugget coverage, and Recall@50 for traditional relevancy.

### v2

@@ 3.4 Retrieval: Oracle & Inference Setting
A RAG evaluation dataset requires questions, answers, and a corpus with documents, which helps support facts in the answer. In this stage, we retrieve a list of highly relevant (unjudged) documents from the document corpus and construct judgment pools. Since, we are constructing an evaluation dataset and we have curated answers for questions, we retrieve documents using two methods: (1) Inference, relying only on the question and automatic approaches, and (2) Oracle, relying on the gold answer or list of nuggets, to pool diverse documents for relevance judgment in the next stage.
Retrieval Settings. We experiment with multiple systems to increase diversity in our judgment pools. First, we experiment with two techniques in the inference setting: (i) GPT-4o Sub-Questions: we decompose the original question and generate synthetic sub-questions with GPT-4o, similar to Rosset et al. [59], concatenated together to retrieve documents, and (ii) GPT-4o Closed Book Answer: we generate a closed-book answer for the original question with GPT-4o, similar to HyDE [17], and use the closed-book answer to retrieve documents. Next, in the oracle setting, we experiment with: (i) Stack Overflow Answer: we use the curated Stack Overflow answer as the question to retrieve documents, and (ii) Stack Overflow Nuggets: we use the list of GPT-4o-generated nuggets (Section 3.3), concatenated as the question to retrieve documents.
Retrieval Models. We experiment with five different code and text-aware retrieval models: (i) BM25, a strong lexical baseline in BEIR [68]. We utilize the default implementation available in Pyserini [40]. (ii) BGE (Gemma-2) [7] a dense retriever model [^4] BGE Gemma-2: https://huggingface.co/BAAI/bge-multilingual-gemma2 fine-tuned on the backbone architecture of Gemma-2 (9B) [57] with an embedding size of 3584 and 8K context length. (iii) E5 Mistral (7B) [75] is a dense retriever model [^5] E5 Mistral 7B: https://huggingface.co/intfloat/e5-mistral-7b-instruct based fine-tuned on the backbone of Mistral 7B [26] with 32 layers and embedding size of 4096. (iv) Voyage-large-2666Voyage-large-2: https://docs.voyageai.com/docs/embeddings is a proprietary and general-purpose embedding model optimized for retrieval quality, with a context length of 16K tokens and embedding size of 1536. (v) Fusion, a hybrid retrieval strategy combining the four individual models, normalizing and summing up the top 100 documents and their scores from each model.
3.5 Nugget-Level Support Assessment with Retrieved Documents
@@
Traditionally, relevance judgments are conducted on selected pools of retrieved documents, i.e., where a human assessor judges the relevance of the question with each provided document. Due to computational costs, recent studies experiment with an LLM judge (instead of a human assessor) for relevance judgments in IR [15, 69, 71, 70, 54]. Questions in existing IR datasets are traditionally short, making it easier to judge document relevance. In contrast, questions in the FreshStack dataset are long and elaborate (between 350–500 tokens in length), containing a mixture of text, code snippets, or outputs, making it challenging to judge question-document relevance directly [13]. For instance, a document may answer a major problem presented in the question, address only part of the question, or contain relevant references and examples, and we need to translate this into a relevance score.
@@
#Q #Docs #GH Avg. N/Q Query Answer Query Answer Rel. Docs/N Rel. Docs/Q
@@
Table 1: FreshStack dataset statistics; Dataset count measures the number of queries, documents, GitHub repositories, and average nuggets per query; Avg. length measures the average text lengths (length distribution in Figure 4); % containing code measures the fraction of queries and answers with code snippets; Relevance judgments measure relevant documents per nugget and per query.
@@
Nugget-level Support. Instead of relying on traditional relevance assessments, we simplify the judgment procedure for GPT-4o and evaluate whether a document supports information (or contains) provided by a nugget. A reminder that a nugget highlights an essential fact of the Stack Overflow question or answer. Judging document relevance at a nugget level is effective as nuggets are factual and short information snippets, reducing the ambiguity often seen during traditional relevance judgments. To reduce computational costs, we evaluate top-k documents (a maximum $k=20$) together with the list of all nuggets for a question in a single inference call ($n+k$). We evaluate support judgment with GPT-4o using a chain-of-thought prompt [78].
4 Dataset Statistics & Evaluation
@@
Completing previous stages, we employ two additional post-processing steps to ensure high-quality question and answer pairs remain in the dataset, sacrificing the overall dataset size. In the first step, we remove unsupported questions, i.e., questions that do not contain even a single relevant document; this removes, on average, 11.8% of the total questions. [^7] Future work may include these questions as they are potentially valuable to answer, and better retrieval systems may be able to find relevant documents. In the next step, we aggressively filter by removing questions containing at least one unsupported nugget, i.e., a nugget not supported by any documents, reducing on average 34.2% of the total questions.
4.1 Dataset Statistics
@@
FreshStack datasets covers five domains for programmers: machine learning, computer vision, backend, front-end, and game development, all listed in Table 1. Stack Overflow topics such as LangChain were introduced in 2023, whereas others, like Laravel or Angular, have questions about the latest versions (e.g., Laravel 10 & 11). Each topic has at least 50 questions, all asked between January 2023 and June 2024 (timeline versus frequency shown in Figure 3). The corpus has at least 25K documents sourced from 4–10 GitHub repositories (repositories listed in Table 7). The questions are even longer than answers (distribution shown in Figure 4), containing 350–500 tokens (computed using GPT-4o tokenizer), and at least 50% of the questions and answers contain code snippets. GPT-4o generates around 3–4 nuggets for each question. Each nugget supports at least 3 relevant documents, resulting in 5–6 relevant documents per question, for all topics.
Method Model LangChain Yolo v7 & v8 Laravel 10 & 11 Angular 16, 17 & 18 Godot4
$\alpha$N@10 C@20 R@50 $\alpha$N@10 C@20 R@50 $\alpha$N@10 C@20 R@50 $\alpha$N@10 C@20 R@50 $\alpha$N@10 C@20 R@50
@@
Table 2: Pooling results by retrieval baselines (including fusion) in inference or oracle settings during FreshStack dataset construction. $\alpha$-N@10 denotes $\alpha$-nDCG@10, C@20 denotes Coverage@20 and R@50 denotes Recall@50. Stack Overflow Answer & Nuggets both rely on the gold answer for retrieval (oracle setting), whereas other methods do not rely on the gold answer for retrieval (inference setting). Overall, we highlight the best result in bold for each setting.
4.2 Retrieval and RAG Evaluation Metrics
@@
IR evaluation traditionally follows the Cranfield paradigm [72], focusing on individual document relevance, independent of other documents. This is used to construct standard test collections, such as BEIR [68] and TREC datasets such as the Deep Learning (DL) track [10, 11, 12]. However, diversity in search [6, 61, 80] penalizes information redundancy within retrieved documents to enrich information content and improve efficiency. Therefore, we evaluate retrieval systems with three metrics with relevance judgments at the nugget-level: $\alpha$-nDCG@10 for diversity and relevance, Coverage@20 for nugget coverage, and Recall@50 for traditional relevancy. For RAG evaluation, we compute the All Strict ($A_{strict}$) metric for nugget-based recall taken from TREC RAG 2024 [51, 52], calculating how many nuggets are supported within a system’s response, where each nugget highlights a different aspect of the answer. Please refer to Appendix C for a detailed overview of each metric.
5 Pooling & Qualitative Evaluation
In this section, we attempt to answer RQ2 by evaluating methods that retrieve documents contributing to the judgment pools. We first evaluate the retrieval baselines during nugget-level support judgment (or sampling pools) with both inference and oracle settings.
In FreshStack, we are constructing a test evaluation dataset. Therefore, we can use the Stack Overflow answer or its variants in constructing judgment pools, as discussed previously in Section 3.4. We pool and sample documents from different systems and techniques, similar to how existing question answering datasets are constructed, such as Natural Questions [33] or XOR-TyDI [4], which assessed the document-level relevance by calculating the answer overlap in the document.
Experimental Settings. We perform retrieval with four techniques and baselines (as explained in Section 3.4) and an ensemble fusion of the top 100 documents, with each model score normalized and summed up. Evaluation metrics include $\alpha$-nDCG@10, Coverage@20, and Recall@50. We use GPT-4o with a temperature setting of 0.1888Separately, we tested temperatures of 0.1 and 0.7, observing an identical downstream retrieval accuracy during FreshStack construction. for both the automatic stages. Nugget generation uses a grading notes prompt with the question and answer, and support assessment uses a chain-of-thought prompt [78], judging up to a maximum of 20 documents simultaneously with a list of nuggets generated for each question. Finally, we sample and judge the top 20 fusion documents from each technique and setting (including the question) to avoid sampling holes, highlighting the importance of document diversity in our judgment pools.
5.1 Document Judgment Pooling Results
We outline the results achieved on document judgment pools during FreshStack construction with techniques from both inference and oracle settings. Key takeaways and findings are discussed below:
Overall Highlights. Table 2 reveals two key findings: (1) Techniques in the oracle setting significantly outperform techniques from the inference setting. We observe that both the Stack Overflow answer and nuggets techniques help pool documents relevant to the question, and (2) fusion outperforms all individual models, highlighting the value of diversity in model choice, aiding in the construction of our judgment pools in niche domains and topics.
Within the inference setting in Table 2, we observe GPT-4o Sub-Questions achieves the best pooling results for four topics (except Yolo v7 & v8), showing that decomposing the question into smaller sub-questions is useful in retrieving relevant documents. Stack Overflow Nuggets achieve the best results (except Laravel 10 & 11) in the oracle setting, showing that breaking down the answer into facts or nuggets is crucial. Amongst the individual models, BM25 achieves the best $\alpha$-nDCG@10 on all topics, asserting the importance of lexical approaches in judgment pool construction.
5.2 Qualitative Analysis
In our work, a crucial component is the automatic construction of nuggets and nugget-level document judgments with GPT-4o. To assess GPT-4o’s accuracy, we calibrate with an expert human assessor (ML researcher) on a subset of LangChain, evaluating the quality of generated nuggets and nugget-document support labels for 60 randomly sampled questions.
Nugget Quality Judgment Quality
Precision 90.1 % Relevant 71.7 %
Recall 96.6 % Partially Relevant 11.7 %
Groundedness 96.4 % Non-Relevant 16.6 %
Table 3: Expert evaluation of GPT-4o nugget quality and nugget-document relevance judgments on 60 sampled queries in LangChain.
5.2.1 Nugget Quality Evaluation
For nugget quality evaluation, we ask the human assessor to answer the following questions ($A$, $B$, and $C$) after reading the Stack Overflow question, answer, and list of nuggets. (1) $A$: Does the nugget produce hallucinated content? requiring a boolean response (2) $B$: Is the information provided in the nugget minor or redundant? also requiring a boolean response. After finishing $A$ and $B$, we ask (3) $C$: How many additional nuggets are required to cover all key ideas, requiring an integer in the response.
Evaluation Metrics. We measure the nugget quality by calculating three metrics, by evaluating: (1) precision (P): whether nuggets generated are accurate, (2) recall (R): whether nuggets cover the key aspects of the answer and (3) groundedness (G): whether nuggets produce non-hallucinated content, i.e., within the scope of the answer. More formally, we define them as follows:
$\textbf{P}=\dfrac{|Nuggets|-\mathrm{sum}(B)}{|Nuggets|},\textbf{R}=\dfrac{|Nuggets|-\mathrm{sum}(B)}{|Nuggets|-\mathrm{sum}(B)+C},\textbf{G}=\dfrac{|Nuggets|-\mathrm{sum}(A)}{|Nuggets|}$ (1)
where $|Nuggets|$ denotes the count of nuggets for a given question.
Experimental Results. As shown in Table 3, nuggets in the topic of LangChain achieve above 90% in precision and 96% in recall and groundedness, indicating GPT-4o can generate high-quality nuggets required in the FreshStack framework. Most nuggets are well-grounded, i.e., do not produce hallucinated content (3.6% error), and cover the key aspects of the answer in terms of recall (3.4% error). Precision errors are higher (9.9% error), showing nuggets may contain either minor or repeated information. Within these errors, the last positioned nugget is not informative in almost 50% of all error cases, and either the first or second positioned nugget in the rest of the error cases.
5.2.2 Relevance Judgment Quality Evaluation
We assess the relevance between nuggets and documents in nugget-level support. Since judging all documents (including negatives) for each nugget is cumbersome, we qualitatively check for precision by evaluating only the relevant pairs. We sample one positive document per question, totaling 60 randomly sampled nugget-document pairs. The annotator labels the relevance on a three-level scale: relevant, partially relevant, or non-relevant.
Model LangChain Yolo v7 & v8 Laravel 10 & 11 Angular 16, 17 & 18 Godot4
$\alpha$N@10 C@20 R@50 $\alpha$N@10 C@20 R@50 $\alpha$N@10 C@20 R@50 $\alpha$N@10 C@20 R@50 $\alpha$N@10 C@20 R@50
Inference Setting: Retrieving documents using only the Stack Overflow (SO) query.
BM25 0.230 0.475 0.261 0.137 0.342 0.337 0.319 0.602 0.441 0.259 0.551 0.340 0.144 0.268 0.200
BM25 + Reranker 0.322 0.587 0.294 0.337 0.590 0.424 0.414 0.729 0.509 0.346 0.647 0.385 0.251 0.407 0.244
BGE (Gemma-2) 0.216 0.548 0.337 0.258 0.547 0.430 0.348 0.699 0.574 0.323 0.571 0.378 0.199 0.479 0.419
BGE (Gemma-2) + Reranker 0.349 0.662 0.387 0.388 0.666 0.459 0.306 0.646 0.571 0.296 0.595 0.387 0.324 0.576 0.471
E5 Mistral (7B) 0.304 0.654 0.393 0.243 0.552 0.394 0.250 0.565 0.470 0.262 0.548 0.368 0.217 0.444 0.359
E5 Mistral (7B) + Reranker 0.385 0.701 0.439 0.364 0.628 0.468 0.305 0.613 0.510 0.306 0.601 0.375 0.315 0.566 0.426
Voyage-large-2 0.246 0.528 0.309 0.270 0.570 0.453 0.345 0.701 0.543 0.304 0.625 0.427 0.282 0.522 0.458
Voyage-large-2 + Reranker 0.345 0.648 0.355 0.418 0.670 0.514 0.302 0.653 0.529 0.300 0.600 0.414 0.342 0.598 0.511
Fusion (4 models) 0.337 0.700 0.477 0.304 0.627 0.534 0.426 0.748 0.646 0.385 0.719 0.532 0.265 0.550 0.505
Fusion (4 models) + Reranker 0.397 0.729 0.501 0.416 0.733 0.592 0.319 0.671 0.614 0.318 0.641 0.488 0.340 0.627 0.545
Best Scores in the Oracle Setting taken from Table 2: Upper Baselines on the FreshStack dataset
SO Answer: Fusion (4 models) 0.484 0.821 0.619 0.546 0.854 0.788 0.564 0.892 0.820 0.470 0.805 0.695 0.449 0.741 0.683
SO Nuggets: Fusion (4 models) 0.519 0.881 0.655 0.601 0.876 0.825 0.566 0.888 0.818 0.544 0.881 0.756 0.476 0.814 0.719
Table 4: Document retrieval results on FreshStack with retrieval and reranker baselines (including fusion). Best scores or upper baselines in the oracle setting are taken from Table 2. The reranker is the Voyage AI rerank-2 model [1] reranking the top 100 documents. If the reranker improves upon the retrieval model, it is highlighted in green else red. We highlight the best result in bold.
6 Main Experiments
In this section, we evaluate retrievers and rerankers on document retrieval and RAG settings on the constructed FreshStack datasets, addressing RQ3 posed in our introduction. All models are evaluated using only the Stack Overflow question to retrieve documents in the inference setting, and do not include any information about the Stack Overflow answer or nuggets, ensuring a fair assessment.
Experimental Settings. We evaluate the same retrieval models used as baselines during pooling in FreshStack: BM25, BGE (Gemma-2), E5 Mistral 7B, Voyage-large-2, and Fusion. In addition, we evaluate the Voyage AI rerank-2 [1] as the reranker with a 16K context length, reranking the top 100 documents retrieved from each first-stage retrieval system and fusion. Metrics used for evaluation are defined in Section 4.2: $\alpha$-nDCG@10, Coverage@20, and Recall@50.
For RAG evaluation, we generate a RAG answer naively with five LLM generators: GPT-4o-mini, GPT-4o, GPT-4.1 (nano, mini), and GPT-4.1. We feed the query and the top 20 retrieved documents concatenated together as context. Next, we evaluate whether the RAG answer supports each nugget generated in Section 3.3, following Pradeep et al. [52], providing three labels: support, partial_support, or no_support. We compute the All Strict ($A_{strict}$) metric for RAG evaluation.
6.1 Document Retrieval Results
Accuracy gap between oracle indicates plenty of headroom. From Table 4, we observe techniques from the oracle setting (using Stack Overflow answers or nuggets) achieve a substantially higher $\alpha$-nDCG@10, Coverage@20, and Recall@50 in contrast to all models, including ensemble fusion and reranking with VoyageAI rerank-2. This highlights the complexity of answering FreshStack questions and demonstrates the headroom for improvement in existing code-mixed retrieval models to decrease the gap between retrieval models at inference and oracle approaches.
Ensemble fusion outperforms individual models. Individual retrieval models demonstrate limited success on the FreshStack dataset; whereas, the ensemble fusion of four retrieval models outperforms each retrieval model across all metrics ($\alpha$-nDCG@10, Coverage@20, and Recall@50) and all five topics, except $\alpha$-nDCG@10 on Godot4. This highlights a crucial point: a compound retrieval system [84], developed as an ensemble of retrieval models or something similar, is required to retrieve documents for niche and challenging topics, at present. However, fusion is inefficient at inference time, as it adds up individual model inferences, requiring alternatives.
Opportunities to improve reranking. When using a weak first-stage retrieval, neural rerankers typically improve document ranking [68], although it has been recently shown that this is not always the case when a strong first-stage retrieval is used [85, 24]. Consistent with these recent observations, reranking provides benefits over BM25 for all topics in the FreshStack dataset. However, for our dense retrievers, reranking provides a clear benefit on some but not all datasets. Specifically, while the reranker enhances $\alpha$-nDCG@10, Coverage@20, and Recall@50 for LangChain, Yolo v7 & v8, and Godot4, it reduces those metrics on Laravel 10 & 11 and Angular 16, 17 & 18 for BGE (Gemma-2), Voyage-large-2 and fusion. We suspect the reranker is better in certain programming languages such as Python, and we keep it as future work to understand the limitations of the neural reranker [24].
Technique Retrieval Generator LangChain Yolo v7 & v8 Laravel 10 & 11 Angular 16, 17 & 18 Godot4
nano mini full nano mini full nano mini full nano mini full nano mini full
Inference Setting: Retrieving documents using only the Stack Overflow query.
Closed No Retrieval GPT-4o 0.395 0.524 0.461 0.591 0.512 0.574 0.486 0.568 0.415 0.518
Book No Retrieval GPT-4.1 0.444 0.517 0.564 0.470 0.663 0.647 0.557 0.646 0.621 0.508 0.604 0.597 0.483 0.616 0.573
Fusion GPT-4o 0.464 0.568 0.477 0.630 0.557 0.635 0.536 0.629 0.452 0.544
StackOverflow Fusion GPT-4.1 0.438 0.578 0.610 0.571 0.649 0.624 0.572 0.668 0.660 0.575 0.670 0.674 0.492 0.573 0.595
Query Fusion + Rerank GPT-4o 0.444 0.587 0.492 0.625 0.545 0.617 0.551 0.620 0.428 0.551
Fusion + Rerank GPT-4.1 0.467 0.594 0.625 0.527 0.679 0.684 0.583 0.657 0.651 0.564 0.663 0.650 0.491 0.578 0.591
Oracle Setting (Upper Baseline): Using the Stack Overflow answer directly or its variants for retrieval of documents within the corpus
StackOverflow Fusion GPT-4o 0.492 0.618 0.559 0.668 0.549 0.656 0.584 0.680 0.477 0.576
Nuggets Fusion GPT-4.1 0.533 0.654 0.651 0.591 0.667 0.667 0.607 0.681 0.696 0.626 0.717 0.709 0.489 0.628 0.668
Table 5: RAG evaluation results measuring nugget recall with All Strict ($A_{strict}$) scores on LLM-generated answer with: GPT-4o-mini and GPT-4o, GPT-4.1 (nano, mini) and GPT-4.1. The knowledge cutoff date for GPT-4o series is October 2023 and GPT-4.1 series is June 2024.
6.2 RAG Evaluation Results
Retrieved context is key for RAG accuracy.
From the evaluation results measuring $A_{strict}$ capturing nugget recall in Table 5, the RAG answer generated in the oracle setting outperforms other techniques except Yolo v7 & v8, indicating that the oracle retrieved context is the key. Consistent across all observations, GPT-4.1 performs better than GPT-4o across all versions and topics, which we suspect is likely due to a more recent knowledge cutoff date (June 2024 versus October 2023). We likely suspect this as a possibility for GPT-4.1 (especially, mini) surprisingly high effectiveness of the closed-book answer. Lastly, the fusion + rerank setting baseline is competitive, outperforming even the oracle setting in Yolo v7 & v8.
7 Conclusion
Appendix A Technical Appendices and Supplementary Material
IR / RAG Benchmarks Niche Complex Dynamic Challenge
Domains Questions Updates Level
CQADupstack [21] No No No Easy
CodeSearchNet [22] No No No Easy
COIR [36] Limited Yes No Moderate
Stack Overflow-QA [36] No Yes No Moderate
CodeRAG-Bench [77] Limited No No Moderate
Neural Code Search [35] No No No Moderate
SWE-Bench [27] No Yes Yes High
FreshStack (ours) Yes Yes Yes High
Table 6: A comparison of existing IR/RAG evaluation benchmarks with FreshStack.
Appendix B Comparison of FreshStack with Existing IR and RAG Benchmarks
Table 6 compares FreshStack against existing code-focused IR or RAG benchmarks. Below, we briefly describe a few advantages of FreshStack over existing RAG benchmarks:
First, the FreshStack framework utilizes user-asked questions and curated answers, making the evaluation challenging. A majority of existing benchmarks are unrealistic, derived from easily retrievable topics and queries, such as Neural Code Search [35], making it easy to retrieve and answer them, rather than being grounded in solving real user problems provided in FreshStack. We are not crafting artificial (or LLM-generated) questions or sampling questions myopically. Second, all answers in FreshStack are supported in real time by information from technical documentation in GitHub repositories. Third, the framework is designed to be general and scalable without modification. Finally, FreshStack is focused on niche domains and recent topics, taking careful measures to mitigate risks with data contamination introduced by LLMs, ensuring that the benchmark is not susceptible to distortion or leaderboard overfitting [63].
Appendix C Retrieval and RAG Evaluation Metrics
C.1 Retrieval Evaluation Metrics